OpenStudy (anonymous):

plz simplify (4b)^0+8

OpenStudy (hoblos):

= 1+8 = 9

OpenStudy (anonymous):

can u explain

OpenStudy (hoblos):

anything raised to the power of 0 equals 1 x^0 = 1 so (4b)^0 = 1

OpenStudy (hoblos):

and 1+8 = 9

OpenStudy (anonymous):

thanks

OpenStudy (hoblos):

my pleasure =)